OpenStudy (anonymous):

What number must you add to complete the square? x2 + 12x = 16

OpenStudy (asnaseer):

in these types of problems, what you are trying to achieve is something of the form:\[(x+a)^2=x^2+2a+a^2\]

OpenStudy (asnaseer):

in your case 2a=12, so a=6 so we have:\[(x+6)^2=x^2+12x+36\]

OpenStudy (anonymous):

If you have x^2 + bx =k then to complete the square you always add (b/2)^2 here b=12 hence you add 36 so that the final equation become x^2+2ax+a^2 so basically you are trying to get a^2.

OpenStudy (asnaseer):

sorry my original statement should have been:\[(x+a)^2=x^2+2ax+a^2\]

OpenStudy (anonymous):

Thanx

OpenStudy (asnaseer):

yw